OpenStudy (anonymous):

If the shadow of a tree is 14 m long and the shadow of a person who is 1.8 m tall is 4 m long, how tall is the tree? Answer to the nearest tenth.

OpenStudy (anonymous):

 HH' is parellall with AC so ABC and BHH' are similar then BH'/BC = BH/AB ... 4/14=1.8/x x=6.3 is tall of the tree